OpenStudy (anonymous):

Explain, in complete sentences, how you would completely factor 15x2 – 25x – 60 by using the grouping method. How would you check your factors for accuracy?

OpenStudy (anonymous):

first, you would factor out 5 to get 5(3x^2-5x-12). You would then factor out whats left in parenthesis.

OpenStudy (anonymous):

yeah i was wrong, @Hope_nicole is right

OpenStudy (anonymous):

thanks @satellite73

OpenStudy (anonymous):

then maybe \[3x^2-5x-12=3x^2-9x+4x-12=3x(x-3)+4(x-3)=(3x+4)(x-3)\]